What volume of 0.10 M H2SO4 must be added to 50ml of a 0.10 M NaOH solution to make a solution in which the molarity of the H2SO4 is 0.050 M? Sir, Please answer fast .

Grade:11

1 Answers

Arun
6 years ago

2 NaOH + H2SO4 → Na2SO4 +2H2O

Thus, first we would have to neutralize the NaOH before H2SO4 can accumulate in the system.

For that, we can use N1V1=N2V2

Normality of H2SO4=2*molarity=0.2N

For NaOH=0.1N

Thus, for neutralization,

V=0.1*50/0.2=25ml

Thus at this point, solution volume=75ml and molarity of H2SO4 is 0M

Now whatever we add will be free H2SO4

Molarity=0.1M of solution added

Thus if V volume more is added the new molarity of the solution will be,

(0.1*V)/(75+V)

We have to find V at which this is 0.05

Thus solving,

(0.1*V)/(75+V)=0.05

Implies V=3.75/0.05=75ml